OpenStudy (yacoub1993):

Find the quotient z1/z2 of the complex numbers. Leave answer in polar form. \[z1 = \frac{ 1 }{ 8 }(\cos \frac{ 2\pi }{ 3 }=i \sin \frac{ 2\pi }{ 3 })\] \[z2= \frac{ 1 }{ 3 }(\cos \frac{ \pi }{ 4 }+i \sin \frac{ \pi }{ 4 })\]

OpenStudy (wolf1728):

Wow - that is beyond my understanding of complex numbers!!

OpenStudy (yacoub1993):

u have no idea what the answer would be

OpenStudy (wolf1728):

Heck no. I don't even understand the question.

OpenStudy (yacoub1993):

lol...hehee anyways thnxs

OpenStudy (wolf1728):

okay u r welcome

OpenStudy (yacoub1993):

@wolf1728 what about this one Polar coordinates of a point are given. Find the rectangular coordinates of the point. (-5, -180°) A. (-5, 0) B. (0, -5) C. (0, 5) D. (5, 0)

OpenStudy (wolf1728):

Rectangular coordinates are properly called Cartesian coordinates. The Cartesian coordinates are (5, 6.123 x 10^-16) I used the calculator here: http://www.engineeringtoolbox.com/converting-cartesian-polar-coordinates-d_1347.html

OpenStudy (wolf1728):

So that makes the answer D

OpenStudy (anonymous):

In polar form \[z=r(\cos \theta + i\sin \theta)=r \ cis\theta=re^{i\theta}\] \[z_1=r_1e^{i\theta}\] \[z_2=r_2e^{i\varphi}\] \[ \frac{z_1}{z_2}=\frac{r_1e^{i\theta}}{r_2e^{i\varphi}}=\frac{r_1}{r_2}e^{i(\theta-\varphi)}\]